You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130220 (20 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+LUCI queue observing, 2013 February 19-20

Observer: E. Sani, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LUCI

Summary

Dome closed at the begining of the night because of strong wind and dewpoint limit.
At 02:00UT it is snowing and the Dome remains closed all night long.
We have problems to insert the calibration unit in LUCI. We call the SA: a power lost during the day caused the lack of energy to the calibration unit. The SA fix it.
Error setting the MOS instrument. We call Dave. A slit mask got stuck, and the SA fix the problem.
The Dome

Data Summary

Project_RUN Objects Filter ExpTime Notes

Details

01:50 We stars with LUCI calib.
  • Problems with the calib unit, we try to pit in IN but nothing happen. We call the SA.

02:24 Run faltFeII.txt for Antoniucci_DDT 0005-0024
  • we increase the exp time for HALO1 0025-34

02:53 Run All_1_200_zJspec.flat for LaFranca_06 0035-0074

03:16 Run M009zCgr32.flat for Zamorani_11
  • Problem with LUCI mask to FPU, error setting the MOS instrument - error committing instrument setup. We call the SA.
  • The error happened because the instrument was not able to remove a longslit mask and insert the PI mask.

05:13 Run M009zCgr32.flat for Zamorani_11 0075-0084
  • with DITx2 0085-0094

05:18 Run M009zCgr32.arc for Zamorani_11 0095-0098
  • with DITx2 0099-0102

05:34 Run All_1_200_zJspec.arc for LaFranca_A9 0103-0114

-- EleanoraSani - 20 Feb 2013
Topic revision: r3 - 20 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback